OpenStudy (anonymous):

Determine if the following statements are true or false A) If Σ|an + bn| converges, then Σ|an| converges and Σ|bn|converges. B) If the terms, a_n, of a series approach zero as n approaches infinity, then the series Σa_n converges.

OpenStudy (anonymous):

A counter-example for B: consider the series \(\displaystyle \sum_{n=1}^\infty\frac{1}{n}.\) This is a divergent series, despite the fact that \(\displaystyle\lim_{n\to\infty}\frac{1}{n}=0\).

OpenStudy (anonymous):

what about A?

OpenStudy (anonymous):

Still thinking about it...